# Neutrino flavor instabilities in a binary neutron star merger remnant: Roles of a long-lived hypermassive neutron star

### v1

@@ Abstract
Understanding the post-merger evolution of binary neutron star merger (BNSM) requires accurate modeling of neutrino transport and microphysics including neutrino flavor conversions. Many previous studies have suggested that fast flavor instability (FFI) and collisional flavor instability (CFI) pervade inner regions of BNSM remnant, and they could impact on fluid dynamics and r-process nucleosynthesis. In this work, we re-examine prospects of occurrences of FFI and CFI using Boltzmann neutrino transport, assuming a frozen fluid background obtained from a numerical relativity simulation of BNSM. We pay special attention to a case involving a long-lived ($>1\,$s) hypermassive neutron star (HMNS). Apart from confirming the claim that these flavor instabilities can occur in BNSM remnants, some new insights are revealed. We identify multiple mechanisms responsible for generating electron neutrino lepton number angular crossings, corresponding to a key indicator of FFI onset, which differ notably from those in black hole (BH) accretion disk systems. We also find that the appearance of positive chemical potential of electron-type neutrinos is a necessary, though not sufficient, condition for FFI, which offers a valuable diagnostic tool for predicting FFI based solely on fluid properties. For CFI, their growth rates are generally lower than FFI, but they can persistently occur in most of the accretion disk up to $\sim 1\,$s. We also find that neglecting contributions of heavy-leptonic neutrinos results in overestimating growth rate and area of unstable regions of CFI. Our result suggests that FFI (CFI) tends to occur transiently (persistently) and locally (widespread in the disk), and FFI is more sensitive to the central compact object (HMNS or BH) than CFI, though more self-consistent simulations with incorporating effects of flavor conversions are needed to confirm these claims.
@@ I Introduction
There is broad agreement in the community that FFI occurs in BNSM remnants. The authors in [24] offer a simple geometric argument of how ELN angular crossings [^1] In general, heavy-leptonic neutrinos and their antipartners have different distributions, implying that XLN contributions need to be taken into account for FFI. On the other hand, XLN is expected to be much less than ELN in BNSM remnant systems. For this reason, all heavy-leptonic neutrinos and their antipartners have been collectively treated even in the modern BNSM simulations (but see [51, 52, 53, 54]). can form. Assuming that there are optically thick regions for both $\nu_{e}$ and $\bar{\nu}_{e}$ in accretion disks (where $\nu_{e}$ and $\bar{\nu}_{e}$ denote electron-type neutrinos and their antipartners, respectively), they suggest that differences in the shapes of these neutrino spheres (or energy spheres) can lead to ELN angular crossings at any spatial positions above the disk. Further investigation in [29], which studies ELN angular crossings based on multi-angle neutrino transport on top of a fluid distribution obtained by realistic BNSM simulations in the early phase ($\sim 5\,$ms after merger), confirm the overall trend. Similarly, the authors in [30, 34, 38] analyze ELN angular crossings based on simulations for black hole (BH) accretion disk system as a model of BNSM remnants. These studies consistently show the presence of ELN crossings, but they also find that underlying mechanisms are more complex than the simple geometric argument [24]. The complexity mirrors findings in CCSN, where multiple mechanisms can give rise to ELN crossings (see, e.g., [42, 55, 56]). Regarding CFI, on the other hand, the authors in [23] show that both non-resonance and resonance CFI can occur in BNSM remnant based on a simulation data of BNSM. However, there are no systematic survey regarding occurrences of CFI. We also note that the study in [23] ignore contributions of heavy-leptonic neutrinos, which potentially leads to false-positive in identifying CFI [57, 45].
@@
In this paper, we present the first systematic study for occurrences of both FFI and CFI based on a long-term simulation of BNSM remnant. We pay special attention to a case with a long-lived HMNS ($>1$ second), which could emerge in coalescenses of relatively low mass two neutron stars ($\lesssim 1.4M_{\odot}$). Such systems are expected to eject larger amounts of neutrinos and baryons from the system compared to short-lived HMNS or prompt formation of BH [58, 59, 60, 61], indicating that flavor instabilities could give large impacts on nucleosynthesis in ejecta. For FFI, we provide a concise interpretation of complex features in generating ELN angular crossings. In addition to this, we propose a necessary (though not sufficient) condition of occurrences for FFI, which will be useful to identify FFI in simulations with approximate (non-multi-angle) neutrino transport. We also discuss some new insights on CFI; for instances their persistency and ubiquity in the accretion disk.
@@ II.1 Remnant model
Figure 1: Profiles of $\nu_{e}$ chemical potential under beta equilibrium ($\mu_{\nu}^{\rm eq}$) in the unit of $[\mathrm{MeV}]$. This can be computed from matter profiles as $\mu_{e}+\mu_{p}-\mu_{n}$. The time displayed in each panel is measured from the timing when we switch to viscous-radiation hydrodynamic simulation. The post-merger time can be obtained by adding $\sim 50\,$ms to them.
@@
In this study, we employ a long-term post-merger BNSM model developed through a hybrid approach of numerical relativity simulations. Assuming equal-mass ($1.35M_{\odot}$) binary neutron stars, the late-inspiral, merger, and early post-merger dynamics (up to $\sim 50\,$ms after merger) were modeled by three dimensional numerical relativity simulations [62]. To efficiently capture the longer-term evolution beyond this point (lasting over 1 second), we switched them to axisymmetric viscous-radiation hydrodynamic simulations [60]. In this study, we employ a model with DD2 equation-of-state (EOS) [63] corresponding to one of the stiff EOSs among those used in our simulations (DD2-135M model in [60]). The HMNS in this model does not enter runaway gravitational collapse, at least not until the end of our simulation ($\sim 1.3\,$s).
In the simulation, neutrino transport is computed based on an energy-integrated formalism as described in [64, 65]. To compute neutrino number density and neutrino-matter interactions (or collision term) within gray neutrino transport schemes, the average energy of neutrinos is estimated by using a prescription in [65]. We incorporate a modern set of weak interactions in BNSM simulations: electron and positron capture of both nucleons and heavy nuclei, electron–positron pair, nucleon–nucleon bremsstrahlung, and plasmon decay; see [66, 67, 68, 69] for details of each weak process.
We note that the present study does not utilize neutrino data directly from the numerical relativity simulation. Instead, we run Boltzmann neutrino transport simulations on top of fixed fluid backgrounds at selected time snapshots ($0,0.2,0.4,0.6,0.8,$ and 1.31.31.3\,1.3s) taken from the viscous-radiation hydrodynamic simulation. These simulations are run until the neutrino radiation field reaches nearly a quasi-steady state. Although this post-processing approach is in some sense redundant, it is essential for accurately capturing angular crossings, which require multi-angle treatment in neutrino transport. In Sec. II.2, we describe some essential information on our numerical scheme for solving Boltzmann equation. It is worth noting that, by avoiding any uncertainties in neutrino angular distributions, we can thoroughly investigate mechanisms for generating ELN angular crossings. These detailed inspections would be difficult for other surrogate prescriptions of crossing search [70, 71, 72, 73, 29].
@@ II.2 Boltzmann neutrino transport
We employ a discrete ordinate (${\rm S_{n}}$) method to solve Boltzmann equation in flat spacetimes under multi-angle, multi-energy, and multi-species treatments [75]. The code was originally designed for the study of CCSN, but it has recently been extended to study neutrino radiation field in BNSM remnants [74]. Thermodynamical quantities and nuclear abundances, which are necessary information on neutrino matter interactions, are determined with DD2 EOS to maintain consistency with numerical relativity simulations. In the collision term, the majority of weak reaction rates are computed based on the formulae in [76], while an extended rate for nucleon–nucleon bremsstrahlung [77] is also adopted. These are the same as those used in our previous study in BNSM remnant [74].
@@
It is important to note that the Boltzmann scheme is different from a neutrino transport module implemented in a CCSN neutrino-radiation hydro code [78, 79, 80, 81]. More specifically, this scheme does not implement the two-energy grid technique [78], updated neutrino-matter interactions [82], and general relativistic effects [83]. However, the Boltzmann solver of [75] offers greater flexibility and extensibility compared to the CCSN code. We also note that the code capability for simulating neutrino transport in BNSM remnant systems has been well tested in [74], exhibiting that it is suited for the purpose of this study.
@@
Assuming axial- and equatorial symmetries in space, we discretize the spatial domain from 00 to 250250250\,250km in radius ($r$) on a spherical polar grid, using $512$ grid points. For the zenith angle ($\theta$), we cover the domain of $20\leq\theta\leq\pi/$ 2 by $96$ grid points. We note that the resolution is higher than our previous study [74]. This is because steep density gradients appear near the surface of HMNS in the late phase, which is mainly due to neutrino cooling. These gradients cause rapid changes in neutrino opacities, especially in regions where neutrinos transition from optically thick to thin medium. To address this, we increase the number of grid points by following a technique similar to that used in our CCSN simulations [84]. The details of the grid structure and our resolution study are found in Appendix B. In neutrino momentum space, the neutrino energy ($\epsilon_{\nu}$) is discritized with 14 grid points in the range of 000\,MeV $<\epsilon_{\nu}<300\,$MeV, and the full solid angle is covered by two angles ($\theta_{\nu}$ for zenith angle and $\phi_{\nu}$ for azimuthal one) with 24 ($0<\theta_{\nu}<\pi$), and 12 ($0<\phi_{\nu}<2\pi$) grid points. For more detail of grid structure in neutrino momentum space, we refer readers to [75].
@@ II.3 Approximate assessments of FFI and CFI
Figure 3: Color maps of growth rates of FFI in the unit of $[\mathrm{s}^{-1}]$. Each panel represents a different time snapshot. Red cross marks in panels of $t=0\,$s and 0.60.60.6\,0.6s represent spatial positions where ELN angular distributions in neutrino momentum space are analyzed to validate our interpretation of the mechanisms for crossing generation; see the text for more details.
@@
where $G_{F}$ denotes Fermi constant and $f_{i}$ represents a distribution function of "$i"$-species neutrinos.
@@ III.1 Global properties
As illustrated in these figures, however, the spatial distributions of these instabilities are highly complex and exhibit significant time dependence. For instances, FFI is absent in the outer region of HMNS ($\gtrsim 20\,$km) during the time interval 0.20.20.2\,0.2s $\lesssim t\lesssim 0.4\,$s; no ELN angular crossings emerge around the pole (z-axis) at any of the time snapshots analyzed in this study; CFI becomes always unstable at the surface of HMNS, while the inner region is always stable; CFI arises in the disk region persistently but they are suppressed at $t\gtrsim 1.3\,$s. These results highlight the need for detailed analyses. In the following, we delve into them for FFI and CFI in Sec. III.2 and III.3, respectively. To facilitate readers’ understanding, we simplify our argument by illustrating the key points graphically for complex mechanisms.
@@ III.2.1 ELN crossings in optically thick regions
Second, one might guess that the impact of FFI originated from these ELN angular crossings is insignificant due to the small size of the spatial region and the shallow depth of the crossings in neutrino momentum space. However, it is premature to make the judgement. While the depth of crossings does affect the growth rate of FFI (see also Eq. (1)), recent studies have shown that even shallow crossings can lead to strong fast flavor conversions (see Sec. III.C in [47]). This suggests that neutrino distributions with flavor conversions can differ substantially from those without FFI. We also note that effects of flavor conversions cannot be confined locally but rather spread by neutrino advections. In fact recent quantum kinetic simulations demonstrate that FFI in small spatial domain can substantially alter global neutrino radiation field [92, 32, 93]. Such pronounced flavor conversions potentially break chemical- and thermal equilibrium with matter, thereby affecting fluid dynamics, chemical compositions, and neutrino emission in BNSM remnant.
@@ III.2.2 ELN crossings induced by $\nu_{e}$ diffusion
Figure 6: Same as Fig. 5 but for $t=0\,$s ($\sim 50\,$ms after merger) time snapshot. The red cross mark corresponds to the spatial position where we display the ELN angular distribution in Fig. 8; see the text for more details.
Figure 7: Schematic picture for ELN crossings induced by $\nu_{e}$ diffusion. The red and blue arrows represent positive ($\nu_{e}>\bar{\nu}_{e}$) and negative ($\nu_{e}<\bar{\nu}_{e}$) ELN, respectively. For convenience in referring to them in the text, the arrows are individually numbered. We note that the letter ”-s” is appended to some arrows to denote scattered neutrinos. The width of each arrow represents the intensity of neutrinos. For this reason, the rays for scattered neutrinos are drawn with thinner lines than those for emitted neutrinos from HMNS. The black filled region represents the high matter density with $\mu_{\nu}^{\rm eq}<0$, corresponds to the major part of HMNS. The yellow one portrays an optically thick region for neutrinos with $\mu_{\nu}^{\rm eq}>0$ (both $\nu_{e}$ and $\bar{\nu}_{e}$ nearly achieve beta equilibrium). The orange filled circles denote neutrons. In this figure, we omit to show protons, since more neutrons than protons exist in the region.
@@
Some important remarks are in order. First, this type of ELN angular crossings does not emerge if the central compact object is a BH due to the absence of strong $\bar{\nu}_{e}$ emission. Second, an accretion torus with $\mu_{\nu}^{\rm eq}>0$ plays a pivotal role in supplying $\nu_{e}$-dominant emission, suggesting that it can be a necessary condition for occurrences of ELN angular crossings. As discussed later, this claim holds in other mechanisms. Third, these ELN angular crossings exhibit strong non-axisymmetric structures in neutrino momentum space. As an example, we provide Fig. 8 displaying a two-dimensional color map of ELN angular distribution at a spatial position where it is highlighted by a cross mark in Fig 6. ELN becomes positive in angular regions of $\cos{\theta_{\nu}}\sim 0$ and $\phi_{\nu}\sim\pi$, which correspond to the neutrino flight directions of Ray 4 and 7 in Fig. 7.

### v2

@@ Abstract
Understanding the post-merger evolution of binary neutron star merger (BNSM) requires accurate modeling of neutrino transport and microphysics including neutrino flavor conversions. Many previous studies have suggested that fast flavor instability (FFI) and collisional flavor instability (CFI) pervade inner regions of BNSM remnant, and they could impact on fluid dynamics and r-process nucleosynthesis. In this work, we re-examine prospects of occurrences of FFI and CFI using Boltzmann neutrino transport, assuming a frozen fluid background obtained from a numerical relativity simulation of BNSM. We pay special attention to a case involving a long-lived (>1>1\,> 1s) hypermassive neutron star (HMNS). Apart from confirming the claim that these flavor instabilities can occur in BNSM remnants, some new insights are revealed. We identify multiple mechanisms responsible for generating electron neutrino lepton number (ELN) angular crossings, corresponding to a key indicator of FFI onset, which differ notably from those in black hole (BH) accretion disk systems. We argue that the appearance of positive chemical potential of electron-type neutrinos plays important roles on generating ELN angular crossings. For CFI, their growth rates are generally lower than FFI, but they can persistently occur in most of the accretion disk up to 1$\sim$ 1\,∼ 1s. We also find that neglecting contributions of heavy-leptonic neutrinos results in overestimating growth rate and area of unstable regions of CFI. Our result suggests that FFI (CFI) tends to occur transiently (persistently) and locally (widespread in the disk), and FFI is more sensitive to the central compact object (HMNS or BH) than CFI, though more self-consistent simulations with incorporating effects of flavor conversions are needed to confirm these claims.
@@ I Introduction
There is broad agreement in the community that FFI occurs in BNSM remnants. The authors in [24] offer a simple geometric argument of how ELN angular crossings [^1] In general, heavy-leptonic neutrinos and their antipartners have different distributions, implying that XLN contributions need to be taken into account for FFI. On the other hand, XLN is expected to be much less than ELN in BNSM remnant systems. For this reason, all heavy-leptonic neutrinos and their antipartners have been collectively treated even in the modern BNSM simulations (but see [51, 52, 53, 54]). can form. Assuming that there are optically thick regions for both $\nu_{e}$ and $\bar{\nu}_{e}$ in accretion disks (where $\nu_{e}$ and $\bar{\nu}_{e}$ denote electron-type neutrinos and their antipartners, respectively), they suggest that differences in the shapes of these neutrino spheres (or energy spheres) can lead to ELN angular crossings at any spatial positions above the disk. Further investigation in [29], which studies ELN angular crossings based on multi-angle neutrino transport on top of a fluid distribution obtained by realistic BNSM simulations in the early phase (5$\sim$ 5\,∼ 5ms after merger), confirm the overall trend. Similarly, the authors in [30, 34, 38] analyze ELN angular crossings based on simulations for black hole (BH) accretion disk system as a model of BNSM remnants. These studies consistently show the presence of ELN crossings, but they also find that underlying mechanisms are more complex than the simple geometric argument [24]. The complexity mirrors findings in CCSN, where multiple mechanisms can give rise to ELN crossings (see, e.g., [42, 55, 56]). Regarding CFI, on the other hand, the authors in [23] show that both non-resonance and resonance CFI can occur in BNSM remnant based on a simulation data of BNSM. However, there are no systematic survey regarding occurrences of CFI. We also note that the study in [23] ignore contributions of heavy-leptonic neutrinos, which potentially leads to false-positive in identifying CFI [57, 45].
@@
In this paper, we present the first systematic study for occurrences of both FFI and CFI based on a long-term simulation of BNSM remnant. We pay special attention to a case with a long-lived HMNS (>1>1> 1 second), which could emerge in coalescenses of relatively low mass two neutron stars ($\lesssim 1.4M_{\odot}$). Such systems are expected to eject larger amounts of neutrinos and baryons from the system compared to short-lived HMNS or prompt formation of BH [58, 59, 60, 61], indicating that flavor instabilities could give large impacts on nucleosynthesis in ejecta. For FFI, we provide a concise interpretation of complex features in generating ELN angular crossings. In addition to this, we propose a necessary (though not sufficient) condition of occurrences for FFI, which will be useful to identify FFI in simulations with approximate (non-multi-angle) neutrino transport. We also discuss some new insights on CFI; for instances their persistency and ubiquity in the accretion disk.
@@ II.1 Remnant model
Figure 1: Profiles of $\nu_{e}$ chemical potential under beta equilibrium ($\mu_{\nu}^{\rm eq}$) in the unit of $[\mathrm{MeV}]$. This can be computed from matter profiles as $\mu_{e}+\mu_{p}-\mu_{n}$. The time displayed in each panel is measured from the timing when we switch to viscous-radiation hydrodynamic simulation. The post-merger time can be obtained by adding 50$\sim$ 50\,∼ 50ms to them.
@@
In this study, we employ a long-term post-merger BNSM model developed through a hybrid approach of numerical relativity simulations. Assuming equal-mass ($1.35M_{\odot}$) binary neutron stars, the late-inspiral, merger, and early post-merger dynamics (up to 50$\sim$ 50\,∼ 50ms after merger) were modeled by three dimensional numerical relativity simulations [62]. To efficiently capture the longer-term evolution beyond this point (lasting over 1 second), we switched them to axisymmetric viscous-radiation hydrodynamic simulations [60]. In this study, we employ a model with DD2 equation-of-state (EOS) [63] corresponding to one of the stiff EOSs among those used in our simulations (DD2-135M model in [60]). The HMNS in this model does not enter runaway gravitational collapse, at least not until the end of our simulation (1.3$\sim$ 1.3\,∼ 1.3s).
In the simulation, neutrino transport is computed based on an energy-integrated formalism as described in [64, 65]. To compute neutrino number density and neutrino-matter interactions (or collision term) within gray neutrino transport schemes, the average energy of neutrinos is estimated by using a prescription in [65]. We incorporate a set of weak interactions in BNSM simulations: electron and positron capture of both nucleons and heavy nuclei, electron–positron pair, nucleon–nucleon bremsstrahlung, and plasmon decay; see [66, 67, 68, 69] for details of each weak process.
We note that the present study does not utilize neutrino data directly from the numerical relativity simulation. Instead, we run Boltzmann neutrino transport simulations on top of fixed fluid backgrounds at selected time snapshots (0,0.2,0.4,0.6,0.8,0,0.2,0.4,0.6,0.8,0 , 0.2 , 0.4 , 0.6 , 0.8 , and 1.31.3\,1.3s) taken from the viscous-radiation hydrodynamic simulation. These simulations are run for 3$\sim$ 3\,∼ 3ms, which is roughly three times longer than the light crossing time of the computational domain ($r\sim 250\,$km). We confirm that the neutrino radiation field reaches nearly a quasi-steady state. It should be cautioned that our analysis may overlook ELN angular crossings caused by short-time (1$\lesssim$ 1\,≲ 1ms) temporal variations in the vicinity of HMNS. Although addressing this issue is beyond the scope of this paper, we plan to extend the present work for more self-consistent treatments with time-dependent Boltzmann neutrino transport (see, e.g., [38]); the results will be reported in a separate paper.
One might think that this post-processing approach is in some sense redundant, but it is essential for accurately capturing angular crossings, which require multi-angle treatment in neutrino transport. In Sec. II.2, we describe some essential information on our numerical scheme for solving Boltzmann equation. It is worth noting that, by avoiding any uncertainties in neutrino angular distributions, we can thoroughly investigate mechanisms for generating ELN angular crossings. These detailed inspections would be difficult for other surrogate prescriptions of crossing search [70, 71, 72, 73, 29].
@@ II.2 Boltzmann neutrino transport
We employ a discrete ordinate (${\rm S_{n}}$) method to solve Boltzmann equation in flat spacetimes [^2] Neglecting general relativistic effects, in particular light-bending effects, causes a systematic error in locating regions occurring ELN angular crossings, because neutrino angular distributions in momentum space hinges on global neutrino transport. As we shall show below, however, ELN angular crossings in optically thick regions, corresponding to strong curvature regimes nearby HMNS, are determined by local properties of matter rather than global ones, whereas general relativistic effects are less important in optically thin region. This suggests that the flat spacetime condition would be reasonable to study qualitative trends in occurrences of ELN angular crossings. under multi-angle, multi-energy, and multi-species treatments [75]. The code was originally designed for the study of CCSN, but it has recently been extended to study neutrino radiation field in BNSM remnants [74]. Thermodynamical quantities and nuclear abundances, which are necessary information on neutrino matter interactions, are determined with DD2 EOS to maintain consistency with numerical relativity simulations. In the collision term, the majority of weak reaction rates are computed based on the formulae in [76], while an extended rate for nucleon–nucleon bremsstrahlung [77] is also adopted. These are the same as those used in our previous study in BNSM remnant [74].
@@
It is important to note that the Boltzmann scheme is different from a neutrino transport module implemented in a CCSN neutrino-radiation hydro code [78, 79, 80, 81]. More specifically, this scheme does not implement the two-energy grid technique [78], updated neutrino-matter interactions [82], and general relativistic effects [83]. Although these improvements have not yet been incorporated, the scheme remains suited to capture qualitative trends of neutrino radiation fields in BNSM remnants. We note that the code capability for simulating neutrino transport in BNSM remnant systems has been well tested in [74].
@@
Assuming axial- and equatorial symmetries in space, we discretize the spatial domain from 0 to 250250\,250km in radius ($r$) on a spherical polar grid, using 512512512 grid points. For the zenith angle ($\theta$), we cover the domain of $0\leq\theta\leq\pi/2$ by 969696 grid points. We note that the resolution is higher than our previous study [74]. This is because steep density gradients appear near the surface of HMNS in the late phase, which is mainly due to neutrino cooling. These gradients cause rapid changes in neutrino opacities, especially in regions where neutrinos transition from optically thick to thin medium. To address this, we increase the number of grid points by following a technique similar to that used in our CCSN simulations [84]. The details of the grid structure and our resolution study are found in Appendix B. In neutrino momentum space, the neutrino energy ($\epsilon_{\nu}$) is discritized with 14 grid points in the range of 00\,MeV $<\epsilon_{\nu}<300\,$MeV, and the full solid angle is covered by two angles ($\theta_{\nu}$ for zenith angle and $\phi_{\nu}$ for azimuthal one [^3] The azimuthal angle in neutrino momentum space is measured from $\boldsymbol{e}_{\theta}$, which corresponds to the spatial coordinate basis of $\theta$; see Fig. 1 in [75].) with 24 ($0<\theta_{\nu}<\pi$), and 12 ($0<\phi_{\nu}<2\pi$) grid points. For more detail of grid structure in neutrino momentum space, we refer readers to [75].
@@ II.3 Approximate assessments of FFI and CFI
Figure 3: Color maps of growth rates of FFI in the unit of $[\mathrm{s}^{-1}]$. Each panel represents a different time snapshot. Red cross marks in panels of $t=0\,$s and 0.60.6\,0.6s represent spatial positions where ELN angular distributions in neutrino momentum space are analyzed to validate our interpretation of the mechanisms for crossing generation; see the text for more details.
@@
[…]
@@ III.1 Global properties
As illustrated in these figures, however, the spatial distributions of these instabilities are highly complex and exhibit significant time dependence. For instances, FFI is absent in the outer region of HMNS (20$\gtrsim$ 20\,≳ 20km) during the time interval 0.20.2\,0.2s $\lesssim t\lesssim 0.4\,$s; no ELN angular crossings emerge around the pole (z-axis) at any of the time snapshots analyzed in this study; CFI becomes always unstable at the surface of HMNS, while the inner region is always stable; CFI arises in the disk region persistently but they are suppressed at $t\gtrsim 1.3\,$s. These results highlight the need for detailed analyses. In the following, we delve into them for FFI and CFI in Sec. III.2 and III.3, respectively. To facilitate readers’ understanding, we simplify our argument by illustrating the key points graphically for complex mechanisms.
@@ III.2.1 ELN crossings in optically thick regions
Second, one might guess that the impact of FFI originated from these ELN angular crossings is insignificant due to the small size of the spatial region, the shallow depth of the crossings in neutrino momentum space, and nearly equal number densities of all species of neutrinos. However, it is premature to make the judgement. While the depth of crossings does affect the growth rate of FFI (see also Eq. (1)), recent studies have shown that even shallow crossings can lead to strong fast flavor conversions (see Sec. III.C in [47]). It is also important to note that, despite similar number densities, neutrino fluxes vary by species due to species-dependent opacities. As shown in previous studies [92, 93, 39, 40], FFC in optically thick regions enhances flavor-integrated fluxes, which facilitate neutrino diffusion and induce a secular cooling effect on fluid dynamics. In addition to this, effects of flavor conversions cannot be confined locally but rather spread by neutrino advections. In fact recent quantum kinetic simulations demonstrate that FFI in small spatial domain can substantially alter global neutrino radiation field [92, 32, 93]. Although more self-consistent studies by neutrino-radiation hydrodynamics incorporating effects of flavor conversions are certainly needed to reach firm conclusion, the FFI might break chemical- and thermal equilibrium with matter, thereby affecting fluid dynamics, chemical compositions, and neutrino emission in BNSM remnant.
@@ III.2.2 ELN crossings induced by $\nu_{e}$ diffusion
Figure 6: Same as Fig. 5 but for $t=0\,$s (50$\sim$ 50\,∼ 50ms after merger) time snapshot. The red cross mark corresponds to the spatial position where we display the ELN angular distribution in Fig. 8; see the text for more details.
@@
Figure 7: Schematic picture for ELN crossings induced by $\nu_{e}$ diffusion. The red and blue arrows represent positive ($\nu_{e}>\bar{\nu}_{e}$) and negative ($\nu_{e}<\bar{\nu}_{e}$) ELN, respectively. For convenience in referring to them in the text, the arrows are individually numbered. We note that the letter ”-s” is appended to some arrows to denote scattered neutrinos. The width of each arrow represents the intensity of neutrinos. For this reason, the rays for scattered neutrinos are drawn with thinner lines than those for emitted neutrinos from HMNS. The black filled region represents the high matter density with $\mu_{\nu}^{\rm eq}<0$, corresponds to the major part of HMNS. The yellow one portrays an optically thick region for neutrinos with $\mu_{\nu}^{\rm eq}>0$ (both $\nu_{e}$ and $\bar{\nu}_{e}$ nearly achieve beta equilibrium). The orange filled circles denote neutrons. In this figure, we omit to show protons, since more neutrons than protons exist in the region. It should also be noted that the neutrino trajectories drawn in this schematic figure are simplified so as for readers to capture the essential trend. We also note that low energy neutrinos (10$\lesssim$ 10\,≲ 10MeV) play an important roles on generating ELN angular crossings. See text for more details.
An important remark should be made here. Although the schematic illustration in Fig. 7 and the accompanying argument are based on a simple ray-tracing description, this is primarily for conceptual clarity. In reality, the situation is complex. Neutrinos experience multiple scatterings as they propagate through the accretion disk, indicating that the neutrinos can not escape freely from the disk but rather diffuse outward. In addition to this, the above discussion omits the importance of energy dependent feature. More specifically, low energy neutrinos can escape from the deeper and dense regions of accretion disk with $\mu_{\nu}^{\rm eq}>0$, whereas high-energy neutrinos decouple at less dense region where $\mu_{\nu}^{\rm eq}<0$. In our model, we confirm that $\nu_{e}$ dominates over $\bar{\nu}_{e}$ for 10$\lesssim$ 10≲ 10MeV, generating positive ELN angular regions. This consideration suggests that the appearance of positive ELN angular crossing depends on the competition between low- and high energy neutrinos. If the excess of $\nu_{e}$ in the low energy is insufficient to overcome the negative ELN in high energy region, crossings are not generated even if the optically thick disk with $\mu_{\nu}^{\rm eq}>0$ is present.
Other important remarks are in order. First, this type of ELN angular crossings does not emerge if the central compact object is a BH due to the absence of strong $\bar{\nu}_{e}$ emission. Second, an accretion torus with $\mu_{\nu}^{\rm eq}>0$ plays a pivotal role in supplying $\nu_{e}$-dominant emission, suggesting that it can be a necessary condition for occurrences of ELN angular crossings. As discussed later, this claim holds in other mechanisms. Third, these ELN angular crossings exhibit strong non-axisymmetric structures in neutrino momentum space. As an example, we provide Fig. 8 displaying a two-dimensional color map of ELN angular distribution at a spatial position where it is highlighted by a cross mark in Fig 6. ELN becomes positive in angular regions of $\cos{\theta_{\nu}}\sim 0$ and $\phi_{\nu}\sim\pi$, which correspond to the neutrino flight directions of Ray 4 and 7 in Fig. 7.